OpenStudy (laddiusmaximus):

Find a positive number x such that the sum of 25x and 1/x is as small as possible.

OpenStudy (anonymous):

\[f(x)=25x+\frac{1}{x}\] \[f'(x)=25-\frac{1}{x^2}\] set = 0 and solve for x to get the critical points

OpenStudy (laddiusmaximus):

25-1/x^(2)=0

Directrix (directrix):

Continuing from above: 25 -1/x^(2) = 0 -1/x^(2) = -25 1/x^2 = 25 And, then ....